Two angles are supplementary if their sum is 180. One angle measures two times the measure of a smaller angle. If x represents the smaller angle and these two angles are supplementary, find the measure of each angle.
Answer by stanbon(75887)

Two angles are supplementary if their sum is 180. One angle measures two times the measure of a smaller angle. If x represents the smaller angle and these two angles are supplementary, find the measure of each angle.
---
smaller = x
supplement = 180-x
----
Equation:
180-x = 2x
180 = 3x
x = 60 (smaller)
180-x = 120 (larger)
======================
Cheers,
Stan H.
=============